OpenStudy (anonymous):

Given: AP=2x squared+17x-11 PB=3x squared+3x-4 AB=90 Find AP.

OpenStudy (mertsj):

We need to know where the points are in relation to one another.

OpenStudy (anonymous):



OpenStudy (mertsj):

So AP+PB =AB 2x^2+17x-11+3x^2=3x-6=90

OpenStudy (mertsj):

5x^2+20x-17=90

OpenStudy (mertsj):

5x^2+20x-15=90 5x^2+20x-105=0 x^2+4x-21=0 (x+7)(x-3)=0 x=-7 or x=3 AP=2(3)^2+17(3)-11 AP=18+51-11=58

OpenStudy (anonymous):

thanks a lot :)

OpenStudy (mertsj):

yw